# Zeroth-Order Constrained Optimization from a Control Perspective via Feedback Linearization

### v1

@@ 7 Conclusions
Large Language Model (LLM) Usage Disclosure
We used a large language model (ChatGPT) to assist with language editing and clarity improvements. All technical content, analysis, and contributions were developed by the authors.
@@ Appendix A Other Related Works on Derivative Free Constrained Optimization
One classical family of approaches is filter methods Audet and Dennis [2004], Echebest et al. [2017], Pourmohamad and Lee [2020], Audet and Dennis [2009], Dzahini et al. [2023], Liuzzi et al. [2010], which are based on pattern search techniques. These methods iteratively reduce the objective function while attempting to decrease constraint violations, often through a progressive barrier function. While conceptually simple, filter methods generally rely on user-specified surrogate functions to generate candidate points and frequently require solving auxiliary subproblems. As a result, they are not easily generalizable to high-dimensional settings.
Another important class of derivative-free optimization techniques is model-based methods Müller and Woodbury [2017], Augustin and Marzouk [2015], Gramacy et al. [2016], which build local surrogate models of the objective and constraints and optimize them iteratively. Such methods can achieve strong performance in low- to medium-dimensional problems, but their reliance on accurate surrogate models makes them sample-intensive and thus less practical in high-dimensional scenarios.
A different perspective is offered by extremum seeking (ES) Ariyur and Krstic [2003], which estimates gradients through deterministic perturbations of the system, typically sinusoidal probing signals, as opposed to random perturbations used in two-point estimators. The estimated gradient is then used to drive the system along a descent flow towards an extremum. ES shares a close connection with zeroth-order optimization: it can be interpreted as the continuous-time counterpart of single-point ZO methods Chen et al. [2022b]. Recent works have begun to extend ES to constrained optimization settings Hazeleger et al. [2022], Chen et al. [2022a], though its relationship to ZO approaches in this regime remains an open direction for future study.
Finally, Bayesian optimization (BO) represents another major branch of derivative-free optimization (see, e.g., Gardner et al. [2014], Hernández-Lobato et al. [2016], Acerbi and Ma [2017]). BO adopts a fundamentally different philosophy: it constructs global probabilistic surrogate models (typically Gaussian processes) and leverages acquisition functions to trade off exploration and exploitation. BO is particularly well-suited for low- to medium-dimensional problems where function evaluations are costly, whereas ZO methods are more appropriate in high-dimensional regimes with relatively inexpensive evaluations.
Appendix B Proof of Theorem 1
Proof of Theorem 1.
@@ Appendix B Proof of Theorem 1
$\quad y_{t+1}-y_{t}=J_{h}(x_{t})(x_{t+1}-x_{t})+\epsilon_{t}$
$=-\eta_{t}J_{h}(x_{t})\left(\widetilde{\nabla}f(x_{t})+\widetilde{J}_{h}(x_{t})^{\top}\lambda_{t}\right)+\epsilon_{t}$
$=-\eta_{t}J_{h}(x_{t})\left(\widetilde{\nabla}f(x_{t})+\widetilde{J}_{h}(x_{t})^{\top}\lambda_{t}^{\star}\right)+\eta_{t}J_{h}(x_{t})\widetilde{J}_{h}(x_{t})^{\top}(\lambda_{t}^{\star}-\lambda_{t})+\epsilon_{t}$
$=-\eta_{t}\left(J_{h}(x_{t})\widetilde{\nabla}f(x_{t})-J_{h}(x_{t})\widetilde{J}(x_{t})^{\top}\left(J_{h}(x_{t})\widetilde{J}(x_{t})^{\top}\right)^{-1}(\widetilde{\nabla}f(x_{t})+Kh(x_{t}))\right)$
$\qquad+\eta_{t}\underbrace{J_{h}(x_{t})\widetilde{J}_{h}(x_{t})^{\top}(\lambda_{t}^{\star}-\lambda_{t})}_{:=\Delta_{t}}+\epsilon_{t}$
$=-\eta_{t}Ky_{t}+\eta_{t}\Delta_{t}+\epsilon_{t}$
$\Longrightarrow$~~ $\quad\|y_{t+1}\|\leq(1-\eta_{t}{\lambda_{\min}}(K))\|y_{t}\|+\eta_{t}\|\Delta_{t}\|+C_{1}\eta_{t}^{2}$
Further, from Lemma 2, we have that
$\|\Delta_{t}\|\leq C_{2}r_{2}^{2},~~\textup{where }C_{2}=nR\left(L_{f}+\frac{64\overline{L}_{h}(nL_{f}\overline{L}_{h}+\|K\|H)}{\underline{L}_{h}^{2}}\right)$
@@
$\|y_{t+1}\|$ $\leq(1-\eta_{t}{\lambda_{\min}}(K))\|y_{t}\|+\eta_{t}C_{2}r_{2}^{2}+C_{1}\eta_{t}^{2}$
$\Longrightarrow\|y_{t}\|$ $\leq\prod_{s=0}^{t-1}(1-\eta_{s}{\lambda_{\min}}(K))\|y_{0}\|+C_{2}r_{2}^{2}\sum_{s=0}^{t-1}\prod_{\tau=s+1}^{t-1}(1-\eta_{\tau}{\lambda_{\min}}(K))\eta_{s}+C_{1}\sum_{s=0}^{t-1}\prod_{\tau=s+1}^{t-1}(1-\eta_{\tau}{\lambda_{\min}}(K))\eta_{s}^{2}$
@@
$\|y_{t}\|$ $\leq(1-\eta{\lambda_{\min}}(K))^{t}\|y_{0}\|+C_{2}r_{2}^{2}\sum_{s=0}^{t-1}(1-\eta{\lambda_{\min}}(K))^{s}\eta+C_{1}\eta^{2}\sum_{s=0}^{t-1}(1-\eta{\lambda_{\min}}(K))^{2}$
$\leq(1-\eta{\lambda_{\min}}(K))^{t}\|y_{0}\|+\frac{C_{2}r_{2}^{2}}{{\lambda_{\min}}(K)}+\frac{C_{1}\eta}{{\lambda_{\min}}(K)}$
@@
$\|y_{t}\|$ $\leq\prod_{s=0}^{t-1}\left(1-\frac{\eta{\lambda_{\min}}(K)}{\sqrt{s+1}}\right)\|y_{0}\|+C_{2}r_{2}^{2}\sum_{s=0}^{t-1}\prod_{\tau=s+1}^{t-1}\left(1-\frac{\eta{\lambda_{\min}}(K)}{\sqrt{\tau+1}}\right)\frac{\eta}{\sqrt{s+1}}$
$\qquad+C_{1}\sum_{s=0}^{t-1}\prod_{\tau=s+1}^{t-1}\left(1-\frac{\eta{\lambda_{\min}}(K)}{\sqrt{\tau+1}}\right)\frac{\eta^{2}}{s}$
@@
$T_{B}\geq 32\left(m\log\left(\frac{192\cdot n\cdot\overline{L}_{h}^{2}}{\underline{L}_{h}^{2}}\right)+\log\left(\frac{T_{G}}{\delta}\right)\right)\sim O\left(m\left(\log(n)+\log\left(\frac{\overline{L}_{h}}{\underline{L}_{h}}\right)\right)+\log\left(\frac{T_{G}}{\delta}\right)\right)$
and
$r_{1}\leq\frac{L_{h}}{8\sqrt{2\overline{L}_{h}R}},\quad r_{2}\leq\frac{L_{h}}{8\sqrt{2n\overline{L}_{h}R}}$
then with probability at least $1-\delta$,
$\|x_{t+1}-x_{t}\|\leq\eta_{t}\left(nL_{f}+\frac{64n\overline{L}_{h}(L_{f}\overline{L}_{h}+\|K\|H)}{\underline{L}_{h}^{2}}\right)$
@@
Further, from Lemma 5, when
$r_{1}\leq\frac{L_{h}}{8\sqrt{2\overline{L}_{h}R}},\quad r_{2}\leq\frac{L_{h}}{8\sqrt{2n\overline{L}_{h}R}},$
we have that
$\sigma_{\min}(G_{h})\geq\frac{\underline{L}_{h}^{2}}{64}.$
Thus
$\|\lambda_{t}\|=\|G_{h}^{-1}(G_{f}-Kh(x_{t}))\|\leq\frac{64(L_{f}\overline{L}_{h}+\|K\|H)}{\underline{L}_{h}^{2}}$
@@
$\|x_{t+1}-x_{t}\|$ $\leq\eta_{t}(\|\widetilde{\nabla}f(x)\|+\|\widetilde{J}_{h}(x_{t})\|\|\lambda_{t}\|)$
$\leq\eta_{t}\left(nL_{f}+\frac{64n\overline{L}_{h}(nL_{f}\overline{L}_{h}+\|K\|H)}{\underline{L}_{h}^{2}}\right),$
which completes the proof. ∎
Lemma 2.
We define an auxiliary variable $\lambda_{t}^{\star}:=-\left(J_{h}(x_{t})\widetilde{J}_{h}(x_{t})^{\top}\right)^{-1}\left(J_{h}(x)\widetilde{\nabla}f(x)-Kh(x)\right)$. Under the conditions as stated in Lemma 1, we have that
@@
$\|J_{h}(x_{t})\widetilde{J}_{h}(x_{t})^{\top}(\lambda_{t}^{\star}-\lambda_{t})\|\leq nR\left(L_{f}+\frac{64\overline{L}_{h}(L_{f}\overline{L}_{h}+\|K\|H)}{\underline{L}_{h}^{2}}\right)r_{2}^{2}$
Proof.
From Taylor’s expansion we have that
$\|G_{h}-J_{h}(x_{t})\widetilde{J}_{h}(x_{t})^{\top}\|\leq n\overline{L}_{h}Rr_{2}^{2},\quad\|G_{f}-J_{h}(x_{t})\widetilde{\nabla}f(x_{t})\|\leq nL_{f}Rr_{2}^{2},\quad\|G_{f}\|\leq nL_{f}\overline{L}_{h}.$
Further from Lemma 5 we have
$\|G_{h}^{-1}\|\leq\frac{64}{\underline{L}_{h}^{2}}.$
And thus
$\quad\|J_{h}(x_{t})\widetilde{J}_{h}(x_{t})^{\top}(\lambda_{t}^{\star}-\lambda_{t})\|$
$=\|J_{h}(x_{t})\widetilde{\nabla}f(x_{t})-Kh(x_{t})+J_{h}(x_{t})\widetilde{J}_{h}(x_{t})^{\top}G_{h}^{-1}(G_{f}-Kh(x_{t}))\|$
$=\|J_{h}(x_{t})\widetilde{\nabla}f(x_{t})-G_{f}+(G_{h}-J_{h}(x_{t})\widetilde{J}_{h}(x_{t})^{\top})G_{h}^{-1}(G_{f}-Kh(x))\|$
$\leq\|J_{h}(x_{t})\widetilde{\nabla}f(x_{t})-G_{f}\|+\|G_{h}-J_{h}(x_{t})\widetilde{J}_{h}(x_{t})^{\top}\|\|G_{h}^{-1}\|(\|G_{f}\|+\|K\|H)$
$\leq n{L}_{f}Rr_{2}^{2}+n\overline{L}_{h}Rr_{2}^{2}\frac{64}{\underline{L}_{h}^{2}}(L_{f}\overline{L}_{h}+\|K\|H)$
@@
$=nR\left(L_{f}+\frac{64\overline{L}_{h}(nL_{f}\overline{L}_{h}+\|K\|H)}{\underline{L}_{h}^{2}}\right)r_{2}^{2}$
Appendix C Proof of Theorem 2
Proof of Theorem 2.
@@ Appendix C Proof of Theorem 2
The proof follows a similar structure as the proof of Theorem 1, with some substantial changes. From Taylor’s expansion and Assumption 3 we know that
$y_{t+1}-y_{t}=J_{h}(x_{t})(x_{t+1}-x_{t})+\epsilon_{t},$
where
@@
$\left(J_{h}(x_{t})\widetilde{J}_{h}(x_{t})^{\top}\right)\lambda_{t}^{\star}+\left(J_{h}(x)\widetilde{\nabla}f(x)-Kh(x)\right)=s^{\star},~~\lambda_{t}^{\star}\geq 0,~~s^{\star}\geq 0,~~(\lambda_{t}^{\star})^{\top}s^{\star}=0$ (20)

### v2

@@ 7 Conclusions
Acknowledgment
The authors acknowledge the use of AI-assisted tools for improving grammar, clarity, and overall writing quality.
@@ Appendix A Other Related Works on Derivative Free Constrained Optimization
One classical family of approaches is filter methods [audet_pattern_2004, echebest_inexact_2017, pourmohamad_statistical_2020, audet_progressive_2009, dzahini_constrained_2023, Luizzi10Sequential], which are based on pattern search techniques. These methods iteratively reduce the objective function while attempting to decrease constraint violations, often through a progressive barrier function. While conceptually simple, filter methods generally rely on user-specified surrogate functions to generate candidate points and frequently require solving auxiliary subproblems. As a result, they are not easily generalizable to high-dimensional settings.
Another important class of derivative-free optimization techniques is model-based methods [muller_gosac_2017, augustin_nowpac_2015, gramacy_modeling_2016], which build local surrogate models of the objective and constraints and optimize them iteratively. Such methods can achieve strong performance in low- to medium-dimensional problems, but their reliance on accurate surrogate models makes them sample-intensive and thus less practical in high-dimensional scenarios.
A different perspective is offered by extremum seeking (ES) [ariyur2003real], which estimates gradients through deterministic perturbations of the system, typically sinusoidal probing signals, as opposed to random perturbations used in two-point estimators. The estimated gradient is then used to drive the system along a descent flow towards an extremum. ES shares a close connection with zeroth-order optimization: it can be interpreted as the continuous-time counterpart of single-point ZO methods [chen2022improve]. Recent works have begun to extend ES to constrained optimization settings [hazeleger_sampled-data_2022, chen_model-free_2022], though its relationship to ZO approaches in this regime remains an open direction for future study.
Finally, Bayesian optimization (BO) represents another major branch of derivative-free optimization (see, e.g., [gardner2014bayesian, hernandez2016general, acerbi2017practical]). BO adopts a fundamentally different philosophy: it constructs global probabilistic surrogate models (typically Gaussian processes) and leverages acquisition functions to trade off exploration and exploitation. BO is particularly well-suited for low- to medium-dimensional problems where function evaluations are costly, whereas ZO methods are more appropriate in high-dimensional regimes with relatively inexpensive evaluations.
Appendix B Proof of Theorem 3.1
Proof B.1 (Proof of Theorem 3.1).
@@
$\textstyle\quad y_{t+1}-y_{t}=J_{h}(x_{t})(x_{t+1}-x_{t})+\epsilon_{t}$
$\textstyle=-\eta_{t}J_{h}(x_{t})\left(\widetilde{\nabla}f(x_{t})+\widetilde{J}_{h}(x_{t})^{\top}\lambda_{t}\right)+\epsilon_{t}$
$\textstyle=-\eta_{t}J_{h}(x_{t})\left(\widetilde{\nabla}f(x_{t})+\widetilde{J}_{h}(x_{t})^{\top}\lambda_{t}^{\star}\right)+\eta_{t}J_{h}(x_{t})\widetilde{J}_{h}(x_{t})^{\top}(\lambda_{t}^{\star}-\lambda_{t})+\epsilon_{t}$
$\textstyle=-\eta_{t}\left(J_{h}(x_{t})\widetilde{\nabla}f(x_{t})-J_{h}(x_{t})\widetilde{J}(x_{t})^{\top}\left(J_{h}(x_{t})\widetilde{J}(x_{t})^{\top}\right)^{-1}(\widetilde{\nabla}f(x_{t})+Kh(x_{t}))\right)$
$\textstyle\qquad+\eta_{t}\underbrace{J_{h}(x_{t})\widetilde{J}_{h}(x_{t})^{\top}(\lambda_{t}^{\star}-\lambda_{t})}_{:=\Delta_{t}}+\epsilon_{t}$
$\textstyle=-\eta_{t}Ky_{t}+\eta_{t}\Delta_{t}+\epsilon_{t}$
$\textstyle\Longrightarrow$ $\textstyle\quad\|y_{t+1}\|\leq(1-\eta_{t}{\lambda_{\min}}(K))\|y_{t}\|+\eta_{t}\|\Delta_{t}\|+C_{1}\eta_{t}^{2}$
Further, from Lemma B.4, we have that
$\textstyle\|\Delta_{t}\|\leq C_{2}r_{2}^{2},~~\textup{where }C_{2}=nR\left(L_{f}+\frac{64\overline{L}_{h}(nL_{f}\overline{L}_{h}+\|K\|H)}{\underline{L}_{h}^{2}}\right)$
@@
$\textstyle\|y_{t+1}\|$ $\textstyle\leq(1-\eta_{t}{\lambda_{\min}}(K))\|y_{t}\|+\eta_{t}C_{2}r_{2}^{2}+C_{1}\eta_{t}^{2}$
$\textstyle\Longrightarrow\|y_{t}\|$ $\textstyle\leq\prod_{s=0}^{t-1}(1-\eta_{s}{\lambda_{\min}}(K))\|y_{0}\|+C_{2}r_{2}^{2}\sum_{s=0}^{t-1}\prod_{\tau=s+1}^{t-1}(1-\eta_{\tau}{\lambda_{\min}}(K))\eta_{s}$
$\textstyle\qquad+C_{1}\sum_{s=0}^{t-1}\prod_{\tau=s+1}^{t-1}(1-\eta_{\tau}{\lambda_{\min}}(K))\eta_{s}^{2}$
@@
$\textstyle\|y_{t}\|$ $\textstyle\!\leq\!(1\!-\!\eta{\lambda_{\min}}(K))^{t}\|y_{0}\|\!+\!C_{2}r_{2}^{2}\sum_{s=0}^{t-1}(1\!-\!\eta{\lambda_{\min}}(K))^{s}\eta\!+\!C_{1}\eta^{2}\sum_{s=0}^{t-1}(1\!-\!\eta{\lambda_{\min}}(K))^{2}$
$\textstyle\leq(1-\eta{\lambda_{\min}}(K))^{t}\|y_{0}\|+\frac{C_{2}r_{2}^{2}}{{\lambda_{\min}}(K)}+\frac{C_{1}\eta}{{\lambda_{\min}}(K)}$
@@
$\textstyle\|y_{t}\|$ $\textstyle\leq\prod_{s=0}^{t-1}\left(1\!-\!\frac{\eta{\lambda_{\min}}(K)}{\sqrt{s+1}}\right)\|y_{0}\|\!+\!C_{2}r_{2}^{2}\sum_{s=0}^{t-1}\prod_{\tau=s+1}^{t-1}\left(1\!-\!\frac{\eta{\lambda_{\min}}(K)}{\sqrt{\tau+1}}\right)\frac{\eta}{\sqrt{s+1}}$
$\textstyle\qquad+C_{1}\sum_{s=0}^{t-1}\prod_{\tau=s+1}^{t-1}\left(1-\frac{\eta{\lambda_{\min}}(K)}{\sqrt{\tau+1}}\right)\frac{\eta^{2}}{s}$
@@
$\textstyle\qquad\qquad\qquad\quad+C_{1}\eta^{2}e^{-\eta{\lambda_{\min}}(K)\sqrt{t}}\sum_{s=1}^{t}e^{\eta{\lambda_{\min}}(K)\sqrt{s}}\frac{1}{s}$
@@
[…]
and $r_{1}\leq\frac{L_{h}}{8\sqrt{2\overline{L}_{h}R}},\quad r_{2}\leq\frac{L_{h}}{8\sqrt{2n\overline{L}_{h}R}}$ then with probability at least $1-\delta$,
$\textstyle\|x_{t+1}-x_{t}\|\leq\eta_{t}\left(nL_{f}+\frac{64n\overline{L}_{h}(L_{f}\overline{L}_{h}+\|K\|H)}{\underline{L}_{h}^{2}}\right)$
@@
Further, from Lemma E.1, when $r_{1}\leq\frac{L_{h}}{8\sqrt{2\overline{L}_{h}R}},r_{2}\leq\frac{L_{h}}{8\sqrt{2n\overline{L}_{h}R}},$ we have that
$\sigma_{\min}(G_{h})\geq\frac{\underline{L}_{h}^{2}}{64}.$ Thus
$\textstyle\|\lambda_{t}\|=\|G_{h}^{-1}(G_{f}-Kh(x_{t}))\|\leq\frac{64(L_{f}\overline{L}_{h}+\|K\|H)}{\underline{L}_{h}^{2}}$
@@
$\textstyle\|x_{t+1}-x_{t}\|$ $\textstyle\leq\eta_{t}(\|\widetilde{\nabla}f(x)\|+\|\widetilde{J}_{h}(x_{t})\|\|\lambda_{t}\|)\leq\eta_{t}\left(nL_{f}+\frac{64n\overline{L}_{h}(nL_{f}\overline{L}_{h}+\|K\|H)}{\underline{L}_{h}^{2}}\right),$
which completes the proof.
Lemma B.4.
We define an auxiliary variable
$\lambda_{t}^{\star}:=-\left(J_{h}(x_{t})\widetilde{J}_{h}(x_{t})^{\top}\right)^{-1}\left(J_{h}(x)\widetilde{\nabla}f(x)-Kh(x)\right)$. Under the conditions as stated in Lemma B.2, we have that
@@
$\textstyle\|J_{h}(x_{t})\widetilde{J}_{h}(x_{t})^{\top}(\lambda_{t}^{\star}-\lambda_{t})\|\leq nR\left(L_{f}+\frac{64\overline{L}_{h}(L_{f}\overline{L}_{h}+\|K\|H)}{\underline{L}_{h}^{2}}\right)r_{2}^{2}$
Proof B.5.
Further from Lemma E.1 we have $\|G_{h}^{-1}\|\leq\frac{64}{\underline{L}_{h}^{2}}.$ And thus
$\textstyle\quad\|J_{h}(x_{t})\widetilde{J}_{h}(x_{t})^{\top}(\lambda_{t}^{\star}-\lambda_{t})\|$
$\textstyle=\|J_{h}(x_{t})\widetilde{\nabla}f(x_{t})-Kh(x_{t})+J_{h}(x_{t})\widetilde{J}_{h}(x_{t})^{\top}G_{h}^{-1}(G_{f}-Kh(x_{t}))\|$
$\textstyle=\|J_{h}(x_{t})\widetilde{\nabla}f(x_{t})-G_{f}+(G_{h}-J_{h}(x_{t})\widetilde{J}_{h}(x_{t})^{\top})G_{h}^{-1}(G_{f}-Kh(x))\|$
$\textstyle\leq\|J_{h}(x_{t})\widetilde{\nabla}f(x_{t})-G_{f}\|+\|G_{h}-J_{h}(x_{t})\widetilde{J}_{h}(x_{t})^{\top}\|\|G_{h}^{-1}\|(\|G_{f}\|+\|K\|H)$
[…]
$\textstyle=nR\left(L_{f}+\frac{64\overline{L}_{h}(nL_{f}\overline{L}_{h}+\|K\|H)}{\underline{L}_{h}^{2}}\right)r_{2}^{2}$
Appendix C Proof of Theorem 3.2
Notations. Throughout this proof, we will use the notation $O(x)$ where $x$ is a positive scalar to denote any variable $\epsilon$ such that $\|\epsilon\|\leq x$. Further, for $A\in\mathbb{R}^{m\times n}$ that has full row rank, we use the notation $P_{A}:=A^{\top}(AA^{\top})^{-1}A$ to denote the orthogonal projector onto the row space of $A$. For a matrix $A\in\mathbb{R}^{m_{1}\times n}$ and $B\in\mathbb{R}^{n\times m_{2}}$. We use $\sigma_{\min}(A|B):=\min_{v=Bu,u\in\mathbb{R}^{m_{2}}}\frac{\|Av\|}{\|v\|}$ to denote the restricted minimum singular value of $A$ in the linear subspace of $B$.
Proof C.1 (Proof of Theorem 3.2).
At timestep $t$ in Algorithm 1, let $U_{t}\!:=[u_{1},u_{2},\dots,u_{T_{B}}]\!\in\!\mathbb{R}^{n\times T_{B}}$, then we have that
$\textstyle\widetilde{\nabla}f(x_{t})=\frac{n}{T_{B}}U_{t}U_{t}^{\top}\nabla f(x_{t})+O(nRr_{1}^{2}),~~\widetilde{J}_{h}(x_{t})=\frac{n}{T_{B}}J_{h}(x_{t})U_{t}U_{t}^{\top}+O(nRr_{1}^{2})$
Further
$\textstyle\quad d_{t}=\widetilde{\nabla}f(x_{t})+\widetilde{J}_{h}(x_{t})^{\top}\lambda_{t}$
[…]
$\textstyle=\frac{n}{T_{B}}U_{t}U_{t}^{\top}\!\!\left(\nabla f(x_{t})\!-\!J_{h}(x_{t})^{\!\top}(J_{h}(x_{t})U_{t}U_{t}^{\top}J_{h}(x_{t})^{\!\top}\!)^{-1}(J_{h}(x_{t})U_{t}U_{t}^{\top}\nabla f(x_{t})\!-\!Kh(x_{t}))\right)$
$\textstyle\quad+\underbrace{O(n\overline{L}_{h}\underline{L}_{h}^{-2}Rr_{1}^{2}+C_{2}r_{2}^{2})}_{\epsilon}$
$\textstyle=\frac{n}{T_{B}}\left(U_{t}(I\!-\!P_{J_{h}U_{t}})U_{t}^{\top}\nabla f(x_{t})+U_{t}U_{t}^{\top}J_{h}(x_{t})^{\top}(J_{h}(x_{t})U_{t}U_{t}^{\top}J_{h}(x_{t})^{\top})^{-1}Kh(x_{t})\right)\!+\!\epsilon$
Given this, we can verify that
$\textstyle J_{h}(x_{t})d_{t}$ $\textstyle=\frac{n}{T_{B}}Kh(x_{t})+\epsilon$
$\textstyle\nabla f(x_{t})^{\top}d_{t}$ $\textstyle=\frac{n}{T_{B}}\left(\|(I-P_{J_{h}U_{t}})U_{t}^{\top}\nabla f(x_{t})\|^{2}\right$.
$\textstyle\left.+\nabla f(x_{t})^{\top}U_{t}U_{t}^{\top}J_{h}(x_{t})^{\top}(J_{h}(x_{t})U_{t}U_{t}^{\top}J_{h}(x_{t})^{\top})^{-1}Kh(x_{t})\right)+\epsilon$
From Assumption 3, we know that $h$ is $R$-smooth and thus
$\textstyle|h(x_{t+1})-h(x_{t})+\eta_{t}J_{h}(x_{t})d_{t}|\leq\frac{M}{2}\|x_{t+1}-x_{t}\|^{2}$
$\textstyle\Longrightarrow$ $\textstyle|h(x_{t+1})-(I-\frac{n}{T_{B}}\eta_{t}K)h(x_{t})|\leq\frac{M}{2}\eta_{t}^{2}\|d_{t}\|^{2}+\epsilon$
$\textstyle\Longrightarrow$ $\textstyle\|h(x_{t+1})\|_{1}-\|h(x_{t})\|_{1}\leq-\eta_{t}\frac{n}{T_{B}}{\lambda_{\min}}(K)\|h(x_{t})\|_{1}+\frac{M}{2}\eta_{t}^{2}\|d_{t}\|^{2}+\epsilon$
Further, from the smoothness of $f$ we have that
$\textstyle\quad f(x_{t+1})-f(x_{t})\leq-\eta_{t}\nabla f(x_{t})^{\top}d_{t}+\frac{M_{f}}{2}\eta_{t}^{2}\|d_{t}\|^{2}$
$\textstyle\leq\eta_{t}\frac{n}{T_{B}}(-\|(I-P_{J_{h}U_{t}})U_{t}^{\top}\nabla f(x_{t})\|^{2}$
$\textstyle\quad+\|\nabla f(x_{t})^{\top}U_{t}U_{t}^{\top}J_{h}(x_{t})^{\top}(J_{h}(x_{t})U_{t}U_{t}^{\top}J_{h}(x_{t})^{\top})^{-1}Kh(x_{t})\|)+\epsilon$
[…]
Thus, by combining the above two inequalities, we have that for
$\tau\geq 64(T_{B}L_{f}+H)\overline{L}_{h}\underline{L}_{h}^{-2}\frac{\|K\|}{{\lambda_{\min}}(K)}$ and $\phi(x):=f(x)+\tau\|h(x)\|_{1}$
[…]
$\textstyle\qquad\qquad\qquad\qquad+\frac{M_{f}+\tau M}{2}\eta_{t}^{2}\|d_{t}\|^{2}+\epsilon$
Further we have that
[…]
hence
$\textstyle\phi(x_{t+1})\!-\!\phi(x_{t})$ $\textstyle\!\leq\!(-\frac{\eta n_{t}}{T_{B}}\!+\!3\eta_{t}^{2}\frac{n^{2}(M_{f}\!+\!\tau M)}{T_{B}})(\|(I\!-\!P_{J_{h}U_{t}})U_{t}^{\top}\nabla f(x_{t})\|^{2}\!\!+\!\|h(x_{t})\|_{1})\!+\!\epsilon$
and thus for $\eta\leq\frac{1}{6n(M_{f}+\tau M)}$ we have
$\textstyle\phi(x_{t+1})-\phi(x_{t})\leq-\frac{n}{2T_{B}}\eta(\|(I-P_{J_{h}U_{t}})U_{t}^{\top}\nabla f(x_{t})\|^{2}+\|h(x_{t})\|_{1})+\epsilon$
$\textstyle=-\frac{n}{2T_{B}}\eta(\|(I-P_{J_{h}U_{t}})U_{t}^{\top}(I-P_{J_{h}})\nabla f(x_{t})\|^{2}+\|h(x_{t})\|_{1})+\epsilon$
From the property of orthogonal projector $P_{J_{h}},P_{J_{h}U_{t}}$ we have
$\textstyle\|(I-P_{J_{h}U_{t}})U_{t}^{\top}(I-P_{J_{h}})\nabla f(x_{t})\|=\inf_{v\in col(J_{h})}\|U_{t}^{\top}((I-P_{J_{h}})\nabla f(x_{t})-v)\|$
$\textstyle\geq\sigma_{\min}(U_{t}^{\top}|[J_{h}(x_{t})^{\!\top}\!\!,\nabla\!f(x_{t})])\inf_{v\in col(J_{h})}\|((I-P_{J_{h}})\nabla f(x_{t})-v)\|$
$\textstyle=\sigma_{\min}(U_{t}^{\top}|[J_{h}(x_{t})^{\!\top}\!\!,\nabla\!f(x_{t})])\|(I-P_{J_{h}})\nabla f(x_{t})\|$
Thus by telescoping we get
$\textstyle\frac{1}{T}\!\sum_{t=1}^{T}\!\!\sigma_{\min}\!(U_{t}^{\top}|[J_{h}(x_{t})^{\!\top}\!\!,\!\nabla\!f(x_{t})])^{2}\|(I\!-\!P_{J_{h}})\nabla f(x_{t})\|^{2}\!\!+\!\|h(x_{t})\|_{1}\!\leq\!\frac{2T_{B}}{\eta n}\!\left(\!\frac{\phi(x_{0})-\phi(T)}{T}\!+\!\epsilon\right)$
$\textstyle\Longrightarrow\!\liminf_{t\to+\infty}\left(\sigma_{\min}(U_{t}^{\top}|[J_{h}(x_{t})^{\!\top}\!\!,\nabla\!f(x_{t})])^{2}\|(I-P_{J_{h}})\nabla f(x_{t})\|^{2}\!+\!\|h(x_{t})\|_{1}\right)\!\leq\!\frac{2T_{B}\epsilon}{\eta n}$
Further, given that the columns of $U_{t}$ is sampled i.i.d. from unit sphere, from Lemma E.5 we know that with a positive probability $(\sigma_{\min}(U_{t}^{\top}|[J_{h}(x_{t})^{\!\top}\!\!,\nabla\!f(x_{t})])^{2}\geq\frac{1}{32}$, thus we have
$\textstyle\liminf_{t\to+\infty}\left(\|(I-P_{J_{h}})\nabla f(x_{t})\|^{2}\!+\!\|h(x_{t})\|_{1}\right)\leq\frac{64T_{B}\epsilon}{\eta n}.$
Appendix D Proof of Theorem 4.1
Proof D.1 (Proof of Theorem 4.1).
@@
The proof follows a similar structure as the proof of Theorem 3.1, with some substantial changes. From Taylor’s expansion and Assumption 3 we know that
$\textstyle y_{t+1}-y_{t}=J_{h}(x_{t})(x_{t+1}-x_{t})+\epsilon_{t},$
@@
(20) $\textstyle\left(\!J_{h}(x_{t})\widetilde{J}_{h}(x_{t})^{\top}\!\right)\lambda_{t}^{\star}\!+\!\left(J_{h}(x)\widetilde{\nabla}f(x)\!-\!Kh(x)\right)\!=\!s^{\star},~\lambda_{t}^{\star}\!\geq\!0,~s^{\star}\!\geq\!0,~(\lambda_{t}^{\star})^{\!\top}s^{\star}\!=\!0.$